OpenStudy (anonymous):

How do I write y=2x-7 in function notation?

OpenStudy (anonymous):

f(x)=2x-7

OpenStudy (anonymous):

I guess f(x) = 2x - 7

OpenStudy (anonymous):

Thats it?

OpenStudy (anonymous):

yup sometimes people like to write y= f(x) = 2x - 7

OpenStudy (anonymous):

Oh wow, I didnt think it was that easy. Thanks!

OpenStudy (anonymous):

I asked the same question when I read yours. Thats it? :P haha. Yes but...it´s just...notation.

OpenStudy (anonymous):

well so long as you understand why y is a function of x so y=f(x) but since y = 2x - 7 therefore f(x)=2x - 7

OpenStudy (anonymous):

but be on the look out cuzz i am pretty sure they will throw a hard one at you where y isn't defined explicitly for x something like 2=2x+y in that case y is not in terms of x so solve for y y=2-2x now you can write f(x)=2-2x just a wee sample.

OpenStudy (anonymous):

Awesome, Thanks! @Marrioncito They took away the ability to give multiple medals :(

OpenStudy (anonymous):

Thats Right

OpenStudy (anonymous):

No problem, I´m not here for the medals. =)

OpenStudy (anonymous):

I just joined 3 days ago. SO i not sure about medals but no one into math is ever rarely in it for medals :P